13. Cooperton Mining just announced it will cut its dividend from $3.99 to $2.38 per share and use the extra funds to expand. Before the​ announcement, Cooperton's dividends were expected to grow at a 3.4%
​rate, and its share price was $48.83. With the planned​ expansion, Cooperton's dividends are expected to grow at a 4.7%
rate. What share price would you expect after the​ announcement? (Assume that the new expansion does not change​ Cooperton's risk.) Is the expansion a good​ investment?
Answer #1

DO(1+G) Valueof Existing stock = = r - G

48.83 = 3.99(1 + 0.034) r – 0.034

48.83r - 1.66022 = 4.12566

r = (4.12566 + 1.66022) / 48.83

r = 11.85%

DO(1 +G)) Valueof New Stock = = T - G

2.38(1+0.047)) 10.1185 -0.047

= $ 34.85

After the announcement the value of share is expected to be $ 34.85

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.